# AgriFM: A Multi-source Temporal Remote Sensing Foundation Model for Crop Mapping

### v2

@@ 2.2 Method
In this section, we introduce our proposed multi-source temporal remote sensing foundation model AgriFM for crop mapping. AgriFM is built based on the modified video swin transformer which can accept multi-source temporal remote sensing images as input and extract spatio-temporal features from various downsampling stages. Upon the pre-training of the foundation model, a versatile decoder is formulated. This decoder is capable of accepting features extracted from various stages of downsampling in the foundation model. It is specifically designed to generate a spectrum of crop mapping results, each correlating to its unique set of labels. The overall structure of the framework is shown in Figure 3.
Figure 3: Structure of foundation model, AgriFM, comprising four stages. The input satellite sequences (MODIS, Landsat, and Sentinel-2) are characterized by specific dimensional parameters: $T$ denotes the temporal length of each sequence (set to 16 frames), while $W$ and $H$ represent the spatial width and height (both fixed at 224 pixels). The number of spectral bands, $C$, varies depending on the data source. The decoder is purposed for the upsampling and fusion of features to yield crop mapping results, each marked by their respective labels.
@@ 2.2.1 Multi-source Temporal Foundation Model
The input data comprises MODIS data with a resolution of 250m, symbolized as $\mathbf{I}_{M}\in\mathcal{R}^{7\times T\times H\times W}$, Landsat data with a resolution of 30m, represented as $\mathbf{I}_{L}\in\mathcal{R}^{6\times T\times H\times W}$, and Sentinel-2 data with a resolution of 10m, denoted as $\mathbf{I}_{S}\in\mathcal{R}^{10\times T\times H\times W}$. While featuring different numbers of spectral channels (7, 6, and 10 bands respectively), all inputs share identical temporal length and spatial dimensions. To process these multi-source satellite data consistently, we implement separate 3D convolutional patch embedding layers that transform each input modality into a unified embedding representation $\mathbf{E}\in\mathcal{R}^{C_{1}\times T_{1}\times H_{1}\times W_{1}}$, where $C_{1}=128$ defines the embedding dimension. $T_{1},H_{1}$ and $W_{1}$ are determined by the patch size parameters $(S_{1},D_{1},D_{1})$ in the embedding module, following the relationships $T_{1}=\lfloor T/S_{1}\rfloor$, $W_{1}=\lfloor W/D_{1}\rfloor$, and $H_{1}=\lfloor H/D_{1}\rfloor$. This spatial-temporal patch decomposition allows the model to maintain the structural relationships in the original data while projecting different satellite inputs into a common feature space suitable for subsequent processing.
@@
The embedding representation is then input into the AgriFM, which is constructed based on a modified video swin transformer (Liu et al., 2022). It is proficient at concurrently extracting spatial and temporal features from the input time-series satellite data. The model comprises four hierarchical stages, each containing $2n$ alternating video swin transformer layers: even-numbered layers utilize window-based multi-head self-attention (W-MSA) for local spatiotemporal modeling within non-overlapping $M\times M\times N$ windows, while odd-numbered layers implement shifted window attention (SW-MSA) with shifts along spatial-temporal axes to establish cross-window connections. This dual-attention mechanism effectively balances computational efficiency with global context modeling capabilities.
A key modification from the original architecture is our synchronized spatiotemporal downsampling operation between stages. Unlike conventional approaches that only reduce spatial dimensions, our patch merging module simultaneously compresses both spatial and temporal resolutions through: (1) spatial downsampling via $D_{i}\times D_{i}$ patch concatenation and linear projection, and (2) temporal reduction through $S_{i}$-frame mean pooling along the time axis. This operation transforms features from stage $i$ as $\mathbf{F}_{i}\in\mathbb{R}^{T_{i}\times H_{i}\times W_{i}\times C_{i}}$ to $\mathbf{X}_{i}\in\mathbb{R}^{T_{i+1}\times H_{i+1}\times W_{i+1}\times C_{i+1}}$, where $T_{i+1}=\lfloor T_{i}/S_{i}\rfloor$, $H_{i+1}=\lfloor H_{i}/D_{i}\rfloor$, and $W_{i+1}=\lfloor W_{i}/D_{i}\rfloor$. The channel dimension expands to $C_{i+1}=2C_{i}$ to preserve information capacity. This progressive spatiotemporal abstraction is particularly crucial for agricultural applications, where long input sequences (typically 16-32 temporal observations) must be distilled into single-frame classification outputs while maintaining discriminative crop growth patterns.
@@ 2.2.2 Supervised Pre-training with Land Cover Fractions
During the pre-training phase, multi-source temporal satellite data are initially fed into the patch embedding module and backbone networks for feature extraction. The input data comprises MODIS data with a resolution of 250m, symbolized as $\mathbf{I_{M}}\in\mathcal{R}^{7\times T\times H\times W}$, Landsat data with a resolution of 30m, represented as $\mathbf{I_{L}}\in\mathcal{R}^{6\times T\times H\times W}$, and Sentinel-2 data with a resolution of 10m, denoted as $\mathbf{I_{S}}\in\mathcal{R}^{10\times T\times H\times W}$. These input data possess varying input channels, but identical time length and size, where $T,H,W$ are 16, 224, 224, respectively.
As previously introduced, we devise multiple patch embedding modules for each different source of satellite data. The embedding for MODIS, Landsat, and Sentinel are respectively symbolized as $\mathbf{E_{M}},\mathbf{E_{L}},\mathbf{E_{S}}$. After embedding, each source of data is transformed into the same dimension, allowing it to be processed by a shared backbone, which is our improved video Swin Transformer. For each patch merging, both temporal and spatial dimensions are downsampled by a factor of two. Subsequently, we use average pooling to convert the output of the final stage, denoted as $\mathbf{X}_{4}$ into a vector $\mathbf{x}$, enabling the estimation of land cover fractions for the input image. We implement a two-layered MLP (multilayer perceptron) network to estimate the land cover fractions $\mathbf{p}$.
@@
where $\hat{\mathbf{p}}_{i}$ and $\mathbf{p}_{i}$ correspond to the estimated fractions and the extracted fractions, respectively.
@@
2.2.3 Versatile Decoder for Unified Crop Mapping
The multi-source temporal foundation model enables the unified crop mapping to process heterogeneous temporal remote sensing data from multiple sources, facilitating feature extraction at different sampling stages. While the foundation model provides powerful feature extraction capabilities, it cannot directly generate crop mapping results. To effectively utilize these multi-scale temporal features, we design a versatile decoder architecture that unifies various crop mapping tasks.
For a given crop mapping task with $K$ input data sources (where $K$ represents any combination of MODIS, Landsat, and Sentinel data), the foundation model first extracts features from each source independently. The decoder reconstructs a single crop map through three upsampling layers, each doubling resolution via upsampling and convolutional layers. Decoder layer $j$ fuses two inputs: 1) upsampled features $U_{j-1}$ from the previous layer (initialized with stage 4 encoder outputs $X_{4}$), and 2) skip-connected features $X_{i}$ from encoder stage $i=4-j$. Each decoder layer contains four key operations: rearrange, interpolate, concat, and feature fusion. Given features $\mathbf{F}_{i}\in\mathbb{R}^{T_{i}\times H_{i}\times W_{i}\times C_{i}}$ from the $i$-th stage of the foundation model, the rearrange operation combines the temporal and channel dimensions to produce $\mathbf{F}_{i}^{{}^{\prime}}\in\mathbb{R}^{H_{i}\times W_{i}\times(T_{i}\times C_{i})}$. The interpolate function then performs upsampling to match the required spatial dimensions. The concat operation merges the upsampled features with the previous decoder layer’s output along the channel dimension, yielding $\mathbf{U}^{c}_{j}\in\mathbb{R}^{H_{i}\times W_{i}\times(K\times C_{i}+C_{i-1})}$. Finally, the feature fusion module, implemented with two convolutional layers and a batch normalization layer, processes the concatenated features to generate the decoder layer output $\mathbf{U}_{j}$.
@@ 2.2.3 Versatile Decoder for Unified Crop Mapping
where $y_{i}$ denotes the ground truth label and $p_{i}$ represents the predicted probability. To address the prevalent class imbalance in crop mapping tasks, we implement a hard sample mining strategy that focuses learning on challenging examples during training.
2.3 Downstream Crop Mapping Tasks and Validation
@@
Since our goal is to construct the foundation model for crop mapping, our validation emphasis lies in ensuring the optimal performance of our method when labels are available. On one hand, we seek to validate the efficacy of our proposed AgriFM in handling a variety of crop mapping tasks, while on the other hand, we aim to demonstrate that it outperforms the existing remote sensing foundation models. To this end, we have selected five representative crop mapping tasks across three regions for validation, including 1) cropland mapping, 2) crop field boundary delineation 3) early-season crops mapping, 4) paddy field mapping, and 5) winter wheat mapping. The satellite data and other detailed information required for validating these tasks are shown in Table 2. For each task, there are corresponding training, validation, and test sets. The spatial distribution of these data is illustrated in Figure 4.
2.3.1 Cropland Mapping, Field Boundary Delineation, and Early-season Crop Mapping
@@
For the tasks of cropland mapping, crop field boundary delineation, and early-season crop mapping in the Auvergne-Rhône-Alpes (ARA) region of France, we derive labels from the EuroCrops dataset (Schneider et al., 2023). EuroCrops is an all-encompassing dataset featuring geo-referenced polygons of agricultural croplands that cover 16 countries within the European Union (EU). It also provides data on specific crop types cultivated across these regions. This dataset allows us to obtain a comprehensive distribution of crops in the region from 2018 to 2020. Using this information, we generate labels for both cropland areas and boundaries, and select 16 crops as labels for the early-season crop mapping task. We utilize Sentinel-2 data from three consecutive years, 2018, 2019, and 2020, for training, validation, and testing respectively. Similar to the pre-training data processing, we aggregate the 10-meter and 20-meter resolution bands of the Sentinel-2 satellite into a unified 10-meter resolution, resulting in data across 10 spectral bands. For each task, we randomly select 2327 images, each measuring $256\times 256$ pixels.
@@
For cropland mapping and field boundary delineation tasks, we utilize the first 32 available scenes from the first six months of each year as model inputs. In contrast, the early-season crop classification task employs a distinct temporal sampling strategy: we select 4 scenes per month to construct a 24-scene time series. This design serves dual purposes: (1) to explicitly validate our model’s capability in handling variable-length temporal inputs, and (2) to incorporate prior knowledge from previous year’s cropping patterns. The prior information, extracted using a ResNet50 (He et al., 2016) backbone, is concatenated with the foundation model’s output features. To maintain computational efficiency while ensuring feature richness, we deliberately limit the input sequence to 24 frames for this specific task. These experiments are designed to assess the effectiveness of our method in addressing high-resolution crop mapping tasks over extended temporal scales.
@@
Figure 4: The study area and dataset detailed information for crop mapping tasks.
Table 2: Detailed information on crop mapping tasks and datasets.
@@
Tasks Region Split Year Sources Resolution Bands Sequences length Images Number Images size
@@
Cropland Mapping Field Boundary delineation ARA, France training 2018 Sentinel-2 10m 10 32 2327 256
@@
Early-season Crop Mapping ARA, France training 2018 Sentinel-2 10m 10 24 2327 256
@@
Paddy Rice Mapping Monsoon Asia training 2019 HLS30 & MODIS 30m &250m 6&7 5&46 3039 224&56
Monsoon Asia validation 2019 HLS30 & MODIS 30m &250m 6&7 5&46 1013 224&56
Monsoon Asia testing 2019 HLS30 & MODIS 30m &250m 6&7 5&46 1013 224&56
@@
2.4 Paddy Rice Mapping in Monsoon Asia

### v3

@@ 2.2 Method
In this section, we introduce our proposed multi-source temporal remote sensing foundation model AgriFM for agriculture mapping. AgriFM is built based on the modified Video Swin Transformer which can accept multi-source temporal remote sensing images as input and extract spatio-temporal features from various downsampling stages. Upon the pre-training of the foundation model, a versatile decoder is formulated. This decoder is capable of accepting features extracted from various stages of downsampling in the foundation model. It is specifically designed to generate a spectrum of mapping results, each correlating to its unique set of labels. The overall structure of the framework is shown in Figure 3.
Figure 3: Structure of foundation model, AgriFM, comprising four stages. The input satellite sequences (MODIS, Landsat-8/9, and Sentinel-2) are characterized by specific dimensional parameters: $T$ denotes the temporal length of each sequence (randonly selected from 3 to 32 frames), while $W$ and $H$ represent the spatial width and height (both fixed at 224 pixels). The number of spectral bands, $C$, varies depending on the data source. The decoder is purposed for the upsampling and fusion of features to yield mapping results, each marked by their respective labels.
@@ 2.2.1 Multi-source Temporal Foundation Model
The input data comprises MODIS data symbolized as $\mathbf{I}_{M}\in\mathcal{R}^{7\times T\times H\times W}$, Landsat-8/9 data represented as $\mathbf{I}_{L}\in\mathcal{R}^{6\times T\times H\times W}$, and Sentinel-2 data denoted as $\mathbf{I}_{S}\in\mathcal{R}^{10\times T\times H\times W}$.
To process these multi-source satellite data consistently, we implement separate 3D convolutional patch embedding modules for each modality. Each module employs a single 3D convolutional layer (Conv3D) that simultaneously performs two operations: dividing the input into non-overlapping spatiotemporal patches and projecting them into an embedding space. The kernel size and stride are set identically to the desired patch size $(S_{1},D_{1},D_{1})$, where $S_{1}$ denotes the temporal patch dimension and $D_{1}$ the spatial dimension. This configuration ensures comprehensive coverage of the input data cube (Channels × Time × Height × Width) without overlap.
For each input patch, the Conv3D operation flattens the pixel values and applies a learned linear projection, outputting a vector of dimension $C_{1}=128$. The resulting sequences of patch embeddings, which are referred to as MODIS tokens $\mathbf{F}_{1}^{M}\in\mathcal{R}^{C_{1}\times T_{1}\times H_{1}\times W_{1}}$, Landsat tokens $\mathbf{F}_{1}^{L}\in\mathcal{R}^{C_{1}\times T_{1}\times H_{1}\times W_{1}}$, and Sentinel tokens $\mathbf{F}_{1}^{S}\in\mathcal{R}^{C_{1}\times T_{1}\times H_{1}\times W_{1}}$, form the standardized input to the AgriFM backbone. The output dimensions are determined by the patch size parameters as $T_{1}=\lfloor T/S_{1}\rfloor$, $W_{1}=\lfloor W/D_{1}\rfloor$, and $H_{1}=\lfloor H/D_{1}\rfloor$.
@@
The backbone of AgriFM is constructed based on a modified Video Swin Transformer (Liu et al., 2022). It is proficient at concurrently extracting spatial and temporal features from the input time-series satellite data. The model comprises four hierarchical stages, each containing $2n$ alternating Video Swin Transformer layers: even-numbered layers utilize window-based multi-head self-attention (W-MSA) for local spatiotemporal modeling within non-overlapping $M\times M\times N$ windows, while odd-numbered layers implement shifted window attention (SW-MSA) with shifts along spatial-temporal axes to establish cross-window connections. This dual-attention mechanism effectively balances computational efficiency with global context modeling capabilities.
Our foundation model backbone is constructed with four hierarchical Video Swin Transformer stages, each designed to progressively extract and refine spatiotemporal features from multi-source satellite data. The first stage begins with the above patch embedding modules that transform raw satellite imageries from different sources into a unified token representation, establishing a common feature space for subsequent processing. The following three stages (2-4) share a consistent structure, each comprising two key components: a patch merging operation that reduces feature map resolution while increasing channel dimension, followed by a series of transformer blocks that perform spatiotemporal feature learning. These transformer blocks maintain the same architecture as the original Video Swin Transformer, employing shifted window attention mechanisms to efficiently capture both local and global dependencies.
This framework is applied consistently across all three satellite data sources (MODIS, Landsat-8/9, and Sentinel-2), with each source processed through identical architectural components. For notational simplicity, we use $\mathbf{F}_{i}$ and $\mathbf{X}_{i}$ without superscripts generically to represent the patch merging outputs and transformer block outputs at stage $i$ for any satellite sources. When referring to specific data source, we use superscript notation: $\mathbf{F}_{i}^{M}$, $\mathbf{F}_{i}^{L}$, $\mathbf{F}_{i}^{S}$ and $\mathbf{X}_{i}^{M}$, $\mathbf{X}_{i}^{L}$, $\mathbf{X}_{i}^{S}$ denote the corresponding features for MODIS, Landsat, and Sentinel-2 data respectively.
Our architectural modification centers on the synchronized spatiotemporal downsampling in patch merging operation between stages, which fundamentally differs from conventional approaches that only reduce spatial dimensions. While conventional approaches typically reduce only spatial dimensions, our patch merging module systematically compresses both spatial and temporal resolutions through a coordinated process.
For spatial downsampling, we employ a comprehensive sampling strategy: along both height and width dimensions, we initiate sampling from starting positions ranging from 0 to $D_{i}-1$ with a consistent stride of $D_{i}$. This method generates exactly $D_{i}\times D_{i}$ distinct sampling patterns, each characterized by unique spatial offsets that collectively ensure complete coverage of the input feature maps. For temporal reduction, we implement mean pooling across $S_{i}$ consecutive frames along the temporal dimension, effectively consolidating temporal information while preserving the essential characteristics of the input sequence. We then concatenate all resulting features along the channel dimension. This concatenated representation subsequently undergoes a linear projection that expands the channel dimension to $C_{i+1}=2C_{i}$, while reducing the spatial dimensions by a factor of $D_{i}$ in both directions and reducing the temporal dimensions by a factor of $S_{i}$.
This operation transforms features from stage $i$, denoted as $\mathbf{F}_{i}\in\mathbb{R}^{T_{i}\times H_{i}\times W_{i}\times C_{i}}$, to $\mathbf{X}_{i}\in\mathbb{R}^{T_{i+1}\times H_{i+1}\times W_{i+1}\times C_{i+1}}$, where the output dimensions are computed as:
$T_{i+1}$ $=\lfloor T_{i}/S_{i}\rfloor,$ (4)
$H_{i+1}$ $=\lfloor H_{i}/D_{i}\rfloor,$
$W_{i+1}$ $=\lfloor W_{i}/D_{i}\rfloor,$
$C_{i+1}$ $=2C_{i}.$
The proposed synchronized spatiotemporal downsampling strategy enables our foundation model to handle dynamically varying input sequence lengths (3-32 frames), making it particularly suitable for agricultural mapping tasks that require flexible temporal modeling across different crop phenological cycles. In the patch embedding module, we set the temporal patch size to 2 for sequences shorter than 16 frames and to 4 for longer sequences, while maintaining a consistent spatial patch size of 4×4. Subsequent patch merging operations (stages 2-4) uniformly apply a temporal downsampling factor of 2. The inherent property of mean pooling ensures graceful handling of edge cases, when the temporal dimension becomes smaller than the downsampling factor, remaining frames are preserved without information loss. This design enables effective processing of various input sequences (3-32 frames) by progressively learning multi-scale temporal patterns through hierarchical stages, capturing phenological characteristics at varying temporal granularities while maintaining computational efficiency.
@@ 2.2.2 Supervised Pre-training with Land Cover Fractions
For pre-training the foundation model, we employ land cover fractions as the supervisory signal to guide the learning of semantically meaningful representations. The feature transformation from the final backbone output to land cover fraction prediction involves a carefully designed computational process. The stage 4 output $\mathbf{X}_{4}\in\mathbb{R}^{T_{4}\times H_{4}\times W_{4}\times C_{4}}$ first undergoes spatiotemporal global average pooling to aggregate information across all temporal and spatial positions:
$\mathbf{x}=\frac{1}{T_{4}\cdot H_{4}\cdot W_{4}}\sum_{t=1}^{T_{4}}\sum_{h=1}^{H_{4}}\sum_{w=1}^{W_{4}}\mathbf{X}_{4}[:,t,h,w].$ (5)
This operation collapses the spatiotemporal dimensions while preserving the channel-wise information, producing a compact feature vector $\mathbf{x}\in\mathbb{R}^{C_{4}}$ that encapsulates the global characteristics of the input sequence. The resulting vector subsequently passes through a single-hidden-layer multilayer perceptron (MLP) that performs the final regression:
$\hat{\mathbf{p}}=\sigma(\mathbf{W}_{2}\cdot\text{ReLU}(\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1})+\mathbf{b}_{2}),$ (6)
where $\mathbf{W}_{1}\in\mathbb{R}^{D\times C_{4}}$ and $\mathbf{W}_{2}\in\mathbb{R}^{K\times D}$ represent the weight matrices of the hidden and output layers respectively, with $D$ denoting the hidden dimension and $K$ the number of land cover categories. The non-linear transformation through ReLU activation enables complex feature interactions, while the final sigmoid function $\sigma$ constrains each element of $\hat{\mathbf{p}}$ to the range [0,1], representing valid fraction estimates. This pre-training framework forces the model to develop representations that directly correspond to physically meaningful land surface properties, establishing a strong foundation for various downstream agricultural monitoring tasks.
@@
where $\hat{\mathbf{p}}_{i}$ and $\mathbf{p}_{i}$ correspond to the estimated fractions and the supervised fractions, respectively.
@@
The mean-teacher framework establishes a symbiotic learning process where the student network learns from both noisy labels and the teacher’s stable predictions, while the teacher gradually integrates the student’s knowledge through exponential moving average (EMA) updates. This mutual refinement, regulated by a high smoothing coefficient ($\alpha$), ensures consistent convergence direction. As pre-training progresses, the teacher’s parameters evolve into a robust temporal ensemble of the student’s weights, effectively filtering label noise while maintaining consensus between both models.
2.2.3 Versatile Decoder for Unified Mapping
The multi-source temporal foundation model enables the unified crop mapping to process heterogeneous temporal remote sensing data from multiple sources, facilitating feature extraction at different sampling stages. While the foundation model provides powerful feature extraction capabilities, it cannot directly generate mapping results. To effectively utilize these multi-scale temporal features, we design a versatile decoder architecture that unifies various agricultrue mapping tasks.
For a given mapping task with $K$ input data sources (where $K$ represents any combination of MODIS, Landsat-8/9, and Sentinel-2 data sources), the foundation model first extracts features from each source independently. The decoder reconstructs a single crop map through three upsampling layers, each doubling resolution via upsampling and convolutional layers. Decoder layer $j$ fuses two inputs: 1) upsampled features $U_{j-1}$ from the previous layer (initialized with stage 4 encoder outputs $X_{4}$), and 2) skip-connected features $X_{i}$ from encoder stage i=4ji=4-j. Each decoder layer contains four key operations: rearrange, interpolate, concat, and feature fusion. Given features $\mathbf{F}_{i}\in\mathbb{R}^{T_{i}\times H_{i}\times W_{i}\times C_{i}}$ from the $i$-th stage of the foundation model, the rearrange operation combines the temporal and channel dimensions to produce $\mathbf{F}_{i}^{{}^{\prime}}\in\mathbb{R}^{H_{i}\times W_{i}\times(T_{i}\times C_{i})}$. The interpolate function then performs upsampling to match the required spatial dimensions. The concat operation merges the upsampled features with the previous decoder layer’s output along the channel dimension, yielding $\mathbf{U}^{c}_{j}\in\mathbb{R}^{H_{i}\times W_{i}\times(K\times C_{i}+C_{i-1})}$. Finally, the feature fusion module, implemented with two convolutional layers and a batch normalization layer, processes the concatenated features to generate the decoder layer output $\mathbf{U}_{j}$.
@@ 2.2.3 Versatile Decoder for Unified Mapping
where $y_{i}$ denotes the ground truth label and $\hat{y}_{i}$ represents the predicted probability. To address the prevalent class imbalance in agriculture mapping tasks, we implement a hard sample mining strategy that focuses learning on challenging examples during training.
2.3 Downstream Mapping Tasks and Validation
@@
Since our goal is to construct the foundation model for agriculture mapping, our validation emphasis lies in ensuring the optimal performance of our method when labels are available. On one hand, we seek to validate the efficacy of our proposed AgriFM in handling a variety of agriculture mapping tasks, while on the other hand, we aim to demonstrate that it outperforms the existing remote sensing foundation models. To this end, we have selected five representative tasks across three regions for validation, including 1) agricultural land mapping, 2) field boundary delineation 3) fine-grained agricultural land use/land cover mapping, 4) paddy field mapping, and 5) winter wheat mapping. The satellite data and other detailed information required for validating these tasks are shown in Table 2. The table provides detailed information on all downstream tasks used for evaluating AgriFM. For each task, we specify the geographic region, data split (training/validation/testing), acquisition year, satellite sources, spatial resolution, spectral bands, temporal length (number of observations per sample), number of images, and spatial size (pixel dimensions). The temporal length indicates the input sequence size for each task. The images column represents the total sample count for each split, while size denotes the spatial dimensions of input patches in pixels. The spatial distribution of these data is illustrated in Figure 4.
2.3.1 Agricultural Land Mapping, Field Boundary Delineation and Agricultural Land Use/Land Cover Mapping
@@
For the tasks of agricultural land mapping, field boundary delineation and agricultural land use/land cover mapping in the Auvergne-Rhône-Alpes (ARA) region of France, we derive labels from the EuroCrops dataset (Schneider et al., 2023). EuroCrops is an all-encompassing dataset featuring geo-referenced polygons of agricultural croplands that cover 16 countries within the European Union (EU). It also provides data on specific crop types cultivated across these regions. This dataset allows us to obtain a comprehensive distribution of crops in the region from 2018 to 2020. Using this information, we generate labels for both agriculture land areas and boundaries, and select 16 agriculture land use/land cover as labels for the agricultural land use/land cover mapping task. We utilize Sentinel-2 data from three consecutive years, 2018, 2019, and 2020, for training, validation, and testing respectively. Similar to the pre-training data processing, we aggregate the 10-meter and 20-meter resolution bands of the Sentinel-2 satellite into a unified 10-meter resolution, resulting in data across 10 spectral bands. For each task, we randomly select 2327 images, each measuring $256\times 256$ pixels.
@@
For agricultural land mapping and field boundary delineation, we utilize the first 32 available scenes from the first six months of each year as model inputs. In contrast, the agricultural land use/land cover mapping task employs a distinct temporal sampling strategy: we select 4 scenes per month to construct a 24-scene time series. This design serves dual purposes: (1) to explicitly validate our model’s capability in handling variable-length temporal inputs, and (2) to incorporate prior knowledge from previous year’s cropping patterns. The prior information, extracted using a ResNet50 (He et al., 2016) backbone, is concatenated with the foundation model’s output features. To maintain computational efficiency while ensuring feature richness, we deliberately limit the input sequence to 24 frames for this specific task. These experiments are designed to assess the effectiveness of our method in addressing high-resolution mapping tasks over extended temporal scales.
@@
Figure 4: The study area and dataset detailed information for downstream mapping tasks.
Table 2: Summary of downstream agricultural mapping tasks and dataset specifications. The table details the geographic region, data split (training/validation/testing), acquisition year, satellite sources, spatial resolution, spectral bands, temporal sequence length, number of sample images, and spatial dimensions for each task. Temporal indicates the number of input frames per sample; Images refers to the total count of labeled instances; Size denotes the spatial dimensions in pixels.
@@
Tasks Region Split Year Sources Resolution Bands Temporal Images Size
@@
Agricultural Land Mapping Field Boundary Delineation ARA, France training 2018 Sentinel-2 10m 10 32 2327 256
@@
Agricultural Land Use /Land Cover Mapping ARA, France training 2018 Sentinel-2 10m 10 24 2327 256
@@
Paddy Rice Mapping Monsoon Asia training 2019 HLS30 30m 6 5 3039 224
Monsoon Asia validation 2019 HLS30 30m 6 5 1013 224
Monsoon Asia testing 2019 HLS30 30m 6 5 1013 224
@@
2.3.2 Paddy Rice Mapping in Monsoon Asia